Calculate the critical angle for

(i) an air-glass surface

(ii) an air-water surface

(iii) a water-glass surface

Draw diagrams in each case illustrating the total reflection of a ray incident on the surface


1
Expert's answer
2021-03-09T15:31:26-0500

Answer

i) the critical angle for an air-glass surface is given

"C'=sin^{-1}(\frac{1}{^a\mu _g}) \\=sin^{-1}(\frac{1}{1}) \\=90^°"

ii) the critical angle for an air-water surface is given

"C''=sin^{-1}(\frac{1}{^a\mu _w}) \\=sin^{-1}(\frac{1}{{1.33}})\\=48°44'"

iii) the critical angle for water-glass surface is given


"C'''=sin^{-1}(\frac{^a\mu _g}{^a\mu _w}) \\=sin^{-1}(\frac{1.55}{{1.33}})\\=62°27'"